Water is being heated in a frictionless piston cylinder device. During the process, 8 Btu of heat is lost to the surroundings, 50,000 lbf-ft of heat is transferred to the container from a burner, and 7 Btu of work is done by the expansion of the vapor. Determine the change in the internal energy for this process.
Answer #1

Total heat added to the container, Q_add = 50000 lbf-ft = (50000*0.00128507) Btu = 64.25 Btu

Total heat lost to the surroundings, Q_lost = 8 Btu

Net heat added to the system , del_Q = Q_add - Q_lost = (64.25 - 8) Btu = 56.25 Btu

Net work done by the vapur, del_W = 7 Btu

According to the first law of thermodynamics,

del_Q = del_U + del_W

change in internal energy, del_U = del_Q - del_W = (56.25 - 7) = 49.25 Btu (ans)

Positive sign in the answer of del_Q indicates that the internal energy of the system increases. (Ans)